# Cost-Efficient Subjective Task Annotation and Modeling through Few-Shot Annotator Adaptation

### v1

@@ B.2 Individual Moral Concern Results
B.3 Overall Results of our Framework
@@
$B=X\%\times\lvert\mathcal{A}\rvert$ $50\%\lvert\mathcal{A}\rvert$ $(\frac{3}{6})$ $66\%\lvert\mathcal{A}\rvert$ $(\frac{4}{6})$ $83\%\lvert\mathcal{A}\rvert$ $(\frac{5}{6})$ $25\%\lvert\mathcal{A}\rvert$ $(\frac{6}{24})$ $50\%\lvert\mathcal{A}\rvert$ $(\frac{12}{24})$ $75\%\lvert\mathcal{A}\rvert$ $(\frac{18}{24})$
@@
Table 6: Overall $F_{1}$ results on Brexit and MFSC dataset for different budgets of annotation $(B)$, with various few shot sampling strategies; mean and standard deviation calculated over repeated runs.
For completeness of our analysis in Section 5.2, Table 6 presents the overall performance ($F^{\text{overall}}_{1}$) for all sampling strategies described in Section 3.
B.4 Effect of the Annotators’ Disagreement
@@
B.5 Hardware Configuration
The experiments were conducted on 4 NVIDIA RTX A6000 gpus with 48GB RAM and overall they were completed in about one week.
@@ B.5 Hardware Configuration
Appendix C Mathematical Symbols
@@
Dataset size #annotators Examples Annotations
Brexit 1120 6
"THE MAJORITY WILL NEVER allow the Mentally Ill Globalists to turn
the world into a SJW and Radical Islam SAFE SPACE #brexit #Trump2016"
"0,0,1,1,0,1"
"A muslim Mayor of London? What!? This PC Sickness has
become a pandemic. England turning into Little Asia. <url>"
"0,0,0,1,1,1"
Not all foreign people who wants to go to the uk have bad intentions.
Improve your law. The #Brexit isn’t gonna help your economy.
"0,0,0,0,0,0"
MFSC 2000 24
’m using the term *ethnicity* in all the meanings. A word doesn’t have an
inner, determined and true meaning; the meaning of a word is steadily
constructing by the speakers of a language. If people in the US use *ethnicity*
as a synonym of *race*, the term *ethnicity* acquires also that meaning,
moreover the anthropological meaning.I guess if they have French flags
celebrating Macron’s victory, they’d likely feel French.",
16 annotators : 0
8 annotators : 1
"More or less. Clinton was regarded as a ""female candidate"", but Le Pen
was a ""far-right candidate"". The left calls for diversity of
ethnicity/gender/etc, but conservatives don’t seem to count. ",
4 annotators : 0
20 annotators:1
"**ing bootlickers, all of them. They WANT a Trump monarchy.
All of them are traitors."
24 annotators: 1
Table 8: Statistics of the datasets used in our experiments, and examples from each alongside their annotations.

### v2

@@ C.1 MFSC (Care label)
C.2 GHC (Hate label)
To ensure the generalizability of our framework, we evaluate it on a larger dataset with an imbalanced number of annotations among annotators. We conducted the experiments using the RoBERTa-Base model due to its superior performance among the base models in our experiments, as well as its resource efficiency.
Gab Hate Corpus (GHC) consists of 27,665 posts from the social network service gab.ai, each annotated by a minimum of three trained annotators, and 18 total annotators. It is coded for hate-based rhetoric and has labels of “assaults on human dignity” or “calls for violence”.
Figure 9: The number of annotated instances by each annotator in GHC dataset
The annotators with less than 1000 annotations were filtered out resulting in 16 annotators. Figure 9 shows the number of annotated instances by each annotator.
Experiments: We replicate the experiment described in Section 4.2 with the same implementation details outlined in Section 4.3. We employ varying budgets of $25\%$, $50\%$, and $75\%$, using the two best-performing sampling methods identified in our experiments ($\mathcal{S}_{bal}$ and $\mathcal{S}_{rand}$), and compare the results to the MTL baseline. The overall results are presented in Table 6. It is evident that our framework consistently outperforms MTL across all numbers of shots, sampling methods, and budget variations. Specifically, with $25\%$ of the budget, our model achieves a gain of $1.6\%$ with $k=64$ and $\mathcal{S}_{rand}$, and with $75\%$ of the budget, our model performs the best, achieving a gain of $2\%$.
$F_{1}^{Overall}\uparrow$ GHC
$25\%$ $50\%$ $75\%$ $100\%$
MTL .417(.004) .433(.007) .442(.013) .451(.006)
$k=16$ $\mathcal{S}_{bal}$45(.004) .46(.002) .464(.003)
$\mathcal{S}_{rand}$455(.008) .469(.005) .468(.003)
$k=32$ $\mathcal{S}_{bal}$456(.002) .459(.001) .464(.003)
$\mathcal{S}_{rand}$461(.003) .472(.001) .468(.001)
$k=64$ $\mathcal{S}_{bal}$458(.004) .461(.002) .466(.003)
$\mathcal{S}_{rand}$467(.003) .474(.002) .468(.002)
$k=128$ $\mathcal{S}_{bal}$466(.001) .466(0) .466(.003)
$\mathcal{S}_{rand}$463(.007) .475(.003) .47(.001)
Table 6: Overall $F_{1}$ scores on GHC dataset, Hate label, with varying annotation budgets (%$(\%B)$.
Impact of the Imbalanced Number of Annotations on Performance Results on the GHC dataset indicate a consistent and significant advantage of our framework, even when applied to larger datasets with imbalanced numbers of annotations across annotators. To further investigate the impact of varying numbers of annotations across annotators on the performance of our framework, we conducted a correlation analysis between each annotator’s performance and their number of annotations. The results revealed no statistically significant correlation between the number of annotations and the overall $F_{1}$ score of an annotator, as indicated by the correlation coefficients for $\mathcal{S}_{rand}$ $(r=-0.17,p=0.25)$ and $\mathcal{S}_{bal}$ $(r=-0.14,p=0.32)$. The plots in Figure 10 illustrate the annotator-level $F_{1}$ scores as the number of annotations of the annotators increases.
Figure 10: $F_{1}$ scores of annotators as the number of annotations increases
Appendix D Additional Details and Results
Here, we present the results of our framework for all values of $k$, with the mean and standard deviations reported for three seeds for the RoBERTa models. Tables 8, 9, and 10 show the results for RoBERTa-Base, RoBERTa-Large, and Llama-3, respectively. The best values are highlighted in bold. As evident, our framework outperforms the baseline across all three models.
D.1 Implementation Details
[…]
Hyperparameter Brexit MFSC
Train Batch Size 16 4
Eval Steps 50 100
Max Length 512 512
Learning Rate 1.2e-04 5e-05
Epochs 10 2
Weight Decay 0.01 0.01
LoRA r 8 4
LoRA Alpha 32 16
LoRA Dropout 0.003 0.1
Table 7: Hyperparameters of the Lora Llama-3 model trained for Brexit and MFSC Datasets
@@ D.1 Implementation Details
$50\%$ $66\%$ $83\%$ $100\%$ $25\%$ $50\%$ $75\%$ $100\%$
@@
Table 8: RoBERTa-Base Overall $F_{1}$ results on Brexit and MFSC datasets for different budgets of annotation $(B)$, with various few shot sampling strategies; mean and standard deviation calculated over repeated runs.
$metric=F_{1}^{overall}\uparrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $100\%$ $25\%$ $50\%$ $75\%$ $100\%$
MTL 0.366(0.123) 0.476(0.026) 0.497(0.012) 0.475(0.012) 0.773(0.002) 0.768(0.007) 0.772(0.004) 0.771(0.004)
$k=16$ $\mathcal{S}_{bal}$ 0.48(0.003) 0.476(0.008) 0.484(0.01) 0.778(0.003) 0.779(0.003) 0.776(0.002)
$\mathcal{S}_{dis}$ 0.475(0.01) 0.471(0.003) 0.487(0.012) 0.787(0.001) 0.787(0.001) 0.779(0.002)
$\mathcal{S}_{mv}$ 0.463(0.027) 0.47(0.003) 0.486(0.018) 0.786(0.001) 0.786(0.002) 0.779(0.002)
$\mathcal{S}_{rand}$ 0.477(0.019) 0.474(0.006) 0.486(0.02) 0.787(0.0) 0.786(0.001) 0.779(0.003)
$k=32$ $\mathcal{S}_{bal}$ 0.486(0.003) 0.479(0.005) 0.486(0.011) 0.777(0.003) 0.778(0.002) 0.776(0.002)
$\mathcal{S}_{dis}$ 0.475(0.006) 0.477(0.009) 0.484(0.014) 0.789(0.002) 0.787(0.001) 0.78(0.002)
$\mathcal{S}_{mv}$ 0.475(0.026) 0.471(0.005) 0.487(0.013) 0.788(0.001) 0.786(0.001) 0.779(0.002)
$\mathcal{S}_{rand}$ 0.485(0.004) 0.473(0.004) 0.485(0.016) 0.787(0.0) 0.787(0.001) 0.779(0.003)
$k=64$ $\mathcal{S}_{bal}$ 0.492(0.006) 0.48(0.002) 0.487(0.009) 0.775(0.004) 0.779(0.002) 0.777(0.001)
$\mathcal{S}_{dis}$ 0.501(0.003) 0.479(0.005) 0.499(0.011) 0.79(0.001) 0.788(0.001) 0.78(0.002)
$\mathcal{S}_{mv}$ 0.479(0.025) 0.474(0.004) 0.487(0.013) 0.79(0.001) 0.787(0.001) 0.78(0.002)
$\mathcal{S}_{rand}$ 0.49(0.008) 0.479(0.007) 0.486(0.004) 0.79(0.002) 0.787(0.001) 0.78(0.003)
$k=128$ $\mathcal{S}_{bal}$ 0.513(0.006) 0.492(0.008) 0.493(0.005) 0.779(0.005) 0.78(0.003) 0.777(0.001)
$\mathcal{S}_{dis}$ 0.506(0.002) 0.486(0.002) 0.497(0.01) 0.791(0.001) 0.79(0.0) 0.781(0.002)
$\mathcal{S}_{mv}$ 0.478(0.031) 0.479(0.005) 0.489(0.014) 0.791(0.001) 0.788(0.0) 0.781(0.002)
$\mathcal{S}_{rand}$ 0.495(0.004) 0.485(0.01) 0.494(0.01) 0.791(0.002) 0.788(0.001) 0.781(0.003)
Table 9: RoBERTa-Large overall Aggregated $F_{1}$ results on BREXIT and MFRC dataset for different %$\%B_{f}$s of annotation, mean and std over 3 runs
$metric=F_{1}^{overall}\uparrow$ Brexit MFSC
$50\%$ $66\%$ $83\%$ $100\%$ $25\%$ $50\%$ $75\%$ $100\%$
MTL 0.335 0.345 0.366 0.351 0.669 0.696 0.715 0.713
$k=16$ $\mathcal{S}_{bal}$ 0.316 0.326 0.355 0.637 0.648 0.692
dis 0.314 0.312 0.353 0.628 0.678 0.683
$\mathcal{S}_{mv}$ 0.318 0.32 0.353 0.66 0.683 0.698
$\mathcal{S}_{rand}$ 0.294 0.32 0.35 0.679 0.693 0.701
$k=32$ $\mathcal{S}_{bal}$ 0.337 0.338 0.363 0.634 0.666 0.696
dis 0.318 0.323 0.353 0.655 0.675 0.699
$\mathcal{S}_{mv}$ 0.329 0.322 0.357 0.681 0.691 0.7
$\mathcal{S}_{rand}$ 0.32 0.329 0.345 0.69 0.693 0.703
$k=64$ $\mathcal{S}_{bal}$ 0.348 0.355 0.373 0.644 0.666 0.69
dis 0.327 0.326 0.351 0.656 0.673 0.691
$\mathcal{S}_{mv}$ 0.359 0.339 0.365 0.685 0.691 0.703
$\mathcal{S}_{rand}$ 0.338 0.332 0.361 0.703 0.706 0.705
$k=128$ $\mathcal{S}_{bal}$ 0.384 0.365 0.379 0.675 0.685 0.704
dis 0.337 0.339 0.357 0.664 0.688 0.698
$\mathcal{S}_{mv}$ 0.365 0.363 0.375 0.698 0.698 0.705
$\mathcal{S}_{rand}$ 0.339 0.343 0.365 0.711 0.713 0.716
Table 10: Llama-3 overall Aggregated $F_{1}$ results on BREXIT and MFRC dataset for different %$\%B_{f}$s of annotation
D.2 Hardware Configuration
The experiments were conducted using four NVIDIA RTX A6000 GPUs, each equipped with 48GB of RAM. The total computation time amounted to approximately 2500 GPU hours. The breakdown of GPU hours for different models is as follows:
Roberta-base experiments: 300 GPU hours
Roberta-large experiments: 600 GPU hours
Llama-3-8B experiments: 1600 GPU hours
D.3 Impact of the Annotators’ Disagreement on Performance
@@
Appendix E Mathematical Symbols